# Xmodel-LM Technical Report

### v1

@@ Abstract
We introduce Xmodel-LM, a compact and efficient 1.1B language model pretrained on over 2 trillion tokens. Trained on our self-built dataset (Xdata), which balances Chinese and English corpora based on downstream task optimization, Xmodel-LM exhibits remarkable performance despite its smaller size. It notably surpasses existing open-source language models of similar scale. Our model checkpoints and code are publicly accessible on GitHub at https://github.com/XiaoduoAILab/XmodelLM.
@@ 2 Pretraining
This chapter details the pretraining process of Xmodel-LM. Firstly, we introduce the sources and composition of our corpus, as well as our preprocessing methods. Secondly, we describe the construction of our customized tokenizer. Finally, we detail the model architecture and training parameter configurations.
@@ 2.1 Training Data
Data processing: For data processing, we are committed to ensuring the quality of the data and reducing its redundancy. We first employ heuristic methods such as paragraph length and punctuation ratio for initial filtering. Subsequently, we utilize a 5-gram Kneser-Ney model based on KenLM Library [Heafield, 2011] to compute text perplexity for further quality filtering. In the next stage, we employed a locality-sensitive hashing method based on SimHash to deduplicate the training data. To balance deduplication quality and efficiency, we implemented a bucketing strategy on the entire dataset, enabling the deduplication process to scale efficiently across large datasets. Finally, we tokenized all datasets using our custom-trained tokenizer, and designed different sampling weights for the datasets based on their characteristics, as shown in Table 1.
@@ 2.2 Tokenizer
We employ the unigram algorithm [Kudo, 2018] for data tokenization, utilizing the implementation provided by Sentence-Piece [Kudo and Richardson, 2018]. In contrast to the extensive vocabularies used in prevailing open-source models, our tokenizer is trained on a mixed corpus of Chinese and English, with a vocabulary size of only 32,000. The comparison of the Xmodel-LM tokenizer with other tokenizers is shown in the Table 2. Despite its smaller size, our tokenizer demonstrates impressive compression rates on test data.
@@ 2.3 Model architecture
RMSNorm. To enhance training stability, we utilize the RMSNorm [Zhang and Sennrich, 2019] normalizing function to normalize the input of each transformer sub-layer, without normalizing the output. Linear layers do not incorporate bias, and word embeddings are not tied.
@@ 2.4 Training
Training is conducted on a single node equipped with 8$\times$×H800 GPUs. To enhance training efficiency and boost Model FLOPS Utilization (MFU), we employ Distributed Data Parallel (DDP) and FlashAttention-V2.
We utilize the cumulative gradient updating method with a mini-batch size of 4 and a gradient accumulation step of 30 per GPU, resulting in a global batch size of 960 with a sequence length of 4096. This setup yields a total token count of 3,932,160 per iteration. Training spans 700,000 iterations, accumulating to a total token count of 2,752,512,000,000.
For optimization, we employ the AdamW optimizer with a maximum learning rate of 6e-4. The learning rate undergoes linear increase from zero to the maximum over the first 2000 updates, followed by annealing to 6e-5 using a cosine schedule. The batch size is set to 4M tokens, with weight decay assigned as 0.1. Additionally, we apply a gradient clipping threshold of 1.0 to regulate the gradient value.
Refer to the training log in the Figure 1, which includes trend graphs showing the training and validation losses as the training token count increases. We used OpenWebText [Gokaslan et al., 2019], which is not included in the training set, as the validation set to calculate the validation loss.
@@ 3 Results
Problem-solving evaluation For exploring the performance of the model beyond common-sense reasoning, we also conducted an evaluation of the model’s problem-solving capability. Specifically, our evaluation comprised:
@@ 5 Conclusions
In summary, our work involve training a 1.1B-sized language model (Xmodel-LM) on a self-built tokenizer and corpus containing both Chinese and English data. Our model demonstrates competitive performance on evaluation datasets compared to models of similar scales. Our efforts contribute to the advancement of knowledge by showcasing the potential of training smaller models with extensive datasets for future applications.

### v2

@@ Abstract
We introduce Xmodel-LM, a compact and efficient 1.1B language model pretrained on around 2 trillion tokens. Trained on our self-built dataset (Xdata), which balances Chinese and English corpora based on downstream task optimization, Xmodel-LM exhibits remarkable performance despite its smaller size. It notably surpasses existing open-source language models of similar scale. Our model checkpoints and code are publicly accessible on GitHub at https://github.com/XiaoduoAILab/XmodelLM.
@@ 2 Pretraining
This chapter details the pretraining process of Xmodel-LM. First, we introduce the sources and composition of our corpus, as well as our preprocessing methods. Second, we describe the construction of our customized tokenizer. Finally, we detail the model architecture and training parameter configurations.
@@ 2.1 Training Data
Data processing: We are committed to ensure the quality of the data and reducing its redundancy. We first employ heuristic methods such as paragraph length and punctuation ratio for initial filtering. Subsequently, we utilize a 5-gram Kneser-Ney model based on KenLM Library [Heafield, 2011] to compute text perplexity for further quality filtering. In the next stage, we employe a locality-sensitive hashing method based on SimHash to deduplicate the training data. To balance deduplication quality and efficiency, we implement a bucketing strategy on the entire dataset, enabling the deduplication process to scale efficiently across large datasets. Finally, we tokenize all datasets using our custom-trained tokenizer, and designed different sampling weights for the datasets based on their characteristics, as shown in Table 1.
@@ 2.2 Tokenizer
We employ the unigram algorithm [Kudo, 2018] for data tokenization, utilizing the implementation provided by Sentence-Piece [Kudo and Richardson, 2018]. In contrast to the extensive vocabularies used in prevailing open-source models, our tokenizer is trained on a mixed corpus of Chinese and English, with a vocabulary size of only 32,000. The comparison of the Xmodel-LM tokenizer with other tokenizers is shown in the Table 2. Despite its small size, our tokenizer demonstrates impressive compression rates on test data.
@@ 2.3 Model architecture
RMSNorm. To enhance training stability, we utilize the RMSNorm [Zhang and Sennrich, 2019] function to normalize the input of each transformer sub-layer, without normalizing the output. Linear layers do not incorporate bias, and word embeddings are not tied.
@@ 2.4 Training
Training is conducted on a single node equipped with 7$\times$×H800 GPUs. To enhance training efficiency and boost Model FLOPS Utilization (MFU), we employ Distributed Data Parallel (DDP) and FlashAttention-V2.
We utilize the cumulative gradient updating method with a mini-batch size of 4 and a gradient accumulation step of 30 per GPU, resulting in a global batch size of 960 with a sequence length of 4096. This setup yields a total token count of 3,440,640 per iteration. Training spans 600,000 iterations, accumulating to a total token count of 2,064,384,000,000.
For optimization, we employ the AdamW optimizer with a maximum learning rate of 6e-4. The learning rate undergoes linear increase from zero to the maximum over the first 2000 updates, followed by annealing to 6e-5 using a cosine schedule. The batch size is set to around 3.5M tokens, with weight decay assigned as 0.1. Additionally, we apply a gradient clipping threshold of 1.0 to regulate the gradient value.
Refer to the training log in the Figure 1, which includes trend graphs showing the training and validation losses as the training token count increases. We use OpenWebText [Gokaslan et al., 2019], which is not included in the training set, as the validation set to calculate the validation loss.
@@ 3 Results
Problem-solving evaluation For exploring the performance of the model beyond common-sense reasoning, we also evaluate the model’s problem-solving capability. Specifically, our evaluation comprised:
@@ 5 Conclusions
In summary, our work involves training a 1.1B-sized language model (Xmodel-LM) on a self-built tokenizer and corpus containing both Chinese and English data. Our model demonstrates competitive performance on evaluation datasets compared to models of similar scales. Our efforts contribute to the advancement of knowledge by showcasing the potential of training smaller models with extensive datasets for future applications.